OpenStudy (anonymous):

the volume of a cube is one hundred forty-four cubic meters. What is the volume of the square based pyramid with the same side length and height? sixteen cubic meters forty-eight cubic meters four hundred thirty-two cubic meters one thousand two hundred ninety-six cubic meters

OpenStudy (ddcamp):

V(cube) = Base Area * height V(pyramid) = (1/3) * Base Area * height If base area and height are the same, the pyramid's volume is a third of the spheres.

OpenStudy (anonymous):

so what should i do to find that formula.

OpenStudy (ddcamp):

That's just the formula for the volume of a pyramid. I know it doesn't really seem intuitive, but you just have to accept it until you prove it in Calc II.

OpenStudy (anonymous):

can you just give me the answer sorry im trying to review for a test and these are my review questions and the lesson did not explain any of it.

OpenStudy (ddcamp):

V = (1/3) * 144 V = 48